Jane Industries manufactures plastic toys. During October, Janes Fabrication Department started work on 10,000 models. DurinPlease provide ALL calculations, I really need to understand this for my test tomorrow
Answer #1

Equivalent unit of material = 10000

Equivalent unit of conversion = (3400*75%+7700+2300*55%) = 11515

Cost per equivalent unit of material = 71000/10000 = 7.10

Cost per equivalent unit of conversion = 240300/11515 = 20.87

Cost of ending WIP = (2300*7.1+1265*20.87) = 42731

So answer is c) $42731